OpenStudy (anonymous):

A polygon has vertices whose coordinates are A(1, 4), B(4, -1), C(-1, -4), and D(-4, 1). The diagonals of this polygon bisect each other. true or false

OpenStudy (anonymous):

@elementwielder

OpenStudy (anonymous):

Also, A circle is inscribed in a square whose vertices have coordinates R(0, 4), S(6, 2), T(4, -4), and U(-2, -2). Find the equation of the circle. (x - 2)² + y² = 68 (x + 2)² + y² = 20 (x - 2)² + y² = 10

OpenStudy (anonymous):

I've done a few types in a row so I don't know which distance formula to use

OpenStudy (anonymous):

Just getting confused

OpenStudy (anonymous):

pythagorean theorum... but 1st find out the distance of each side... i recommend drawing the graph on a paper then making right triangles and using the pythagorean thm to solve for all lines

OpenStudy (anonymous):

I know this is unreasonable but would you mind please doing it? I really need to finish and i have a deadline and I have been doing doing this math nonstop for hour.My mind is fried

OpenStudy (anonymous):

hours

OpenStudy (anonymous):

What's the equation

OpenStudy (anonymous):

you dont need an equation for these... you just need to count the differences

OpenStudy (anonymous):

Of the lines you drew?

OpenStudy (anonymous):

yes so the horizontal line just count between the x-values of point C and B and count the differences of the points using the y-values for the vertical line

OpenStudy (anonymous):

3 and 3? Is that what you mean?

OpenStudy (anonymous):

I'm literally brain dead right now...

OpenStudy (anonymous):

ok ill solve for the sides that i just asked you to solve if you can find the line ill point to using the pythagorean theorum

OpenStudy (anonymous):

Okay thank you. So what it line am I looking for, the one with the black arrow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And I find that by using the coordinates of C and B right? and okay

OpenStudy (anonymous):